# Fine-Grained Reward Optimization for Machine Translation using Error Severity Mappings

### v2

@@ 1 Introduction
Neural machine translation (NMT) Kalchbrenner and Blunsom (2013); Sutskever et al. (2014); Cho et al. (2014), a leading approach within MT, leverages neural networks to automate language translation and has driven significant improvements in translation quality. However, most NMT systems are predominantly trained using maximum likelihood estimation (MLE). MLE-based training focuses on maximizing the probability of next-word predictions given a partial reference. This often leads to a critical problem known as exposure bias – the model uses ground-truth prefix tokens during training, but during inference it relies on its previous predictions Bengio et al. (2015); Ranzato et al. (2015); Wiseman and Rush (2016). This can cause errors to propagate through the generated sequence, severely degrading the translation quality. Furthermore, it tends to produce translations that lack global coherence and adequacy as the model does not sufficiently consider the context of entire sentences or the overarching meaning. This has spurred interest in using alternative approaches that leverage RL methods for training NMT systems.
@@
RL-based approaches use explicit reward models to evaluate the outputs generated by the NMT system, assigning scores to generated hypotheses to guide the learning process. However, most prior research Ranzato et al. (2015); Wu et al. (2016); Bahdanau et al. (2016); Nguyen et al. (2017); Wu et al. (2017); Kreutzer et al. (2018a, b); Kiegeland and Kreutzer (2021) predominantly relies on sentence-level feedback and often struggles with reward sparsity, particularly for long-form text generation: sentence-level rewards fail to capture specific issues within a translation, making it difficult for the model to learn from negative reward signals. As shown in Figure 1, two translations corresponding to different source texts of varying length receive the same sentence-level quality score of 70, yet differ significantly in the nature and impact of the errors: The first translation has several minor errors scattered throughout the text, while the latter has major errors that could potentially hinder the understanding of the original content. This suggests that learning can be more effective if feedback is provided at a fine-grained level, including precise identification of the nature of errors.
@@
Recent advancements in automated MT evaluation metrics that generate fine-grained error span predictions, such as xCOMET Guerreiro et al. (2024), MetricX Juraska et al. (2023), AutoMQM Fernandes et al. (2023), and MaTESE Perrella et al. (2022), have shown promise in improving alignment with human translation quality judgments. These metrics directly predict token-level error severity (no error, minor, major, or critical) and optionally sentence-level quality assessments, or prompt large language models to identify error types (e.g. mistranslation, omission) and severities based on the Multidimensional Quality Metrics (MQM) framework (Lommel et al., 2014).
@@
Despite the potential of the severity-based metrics to improve translation quality, their application in MT training via RL methods remains relatively underexplored, since it presents several challenges: (i) the feedback, albeit informative and frequent, can be noisy, and (ii) determining the appropriate reward assignments for different severity levels to ensure effective and stable learning is not straightforward. In this regard, our research aims to answer the following questions:
@@
Can automatic fine-grained MT metrics be effectively used to provide accurate, detailed, human-aligned feedback to reduce reward sparsity?
@@
By integrating fine-grained reward signals into NMT training, we demonstrate significant improvements in translation quality and overcome the challenges of exposure bias, reward sparsity, and stability of RL training, paving the way for more reliable and accurate MT systems.
@@ 2 Background
In the context of MT, we can model the translation process as a Markov Decision Process (MDP) Puterman (1990), defined by the tuple with a finite vocabulary . The state space consists of all possible sequences of tokens up to the current time step, which includes the input sequence in the source language as well as the target language tokens generated so far. Initially, the state corresponds to the input sentence in the source language, , where each token . At each time step , the state represents the sequence of tokens generated up to that point, which can be expressed as:
The agent selects an action , which is a token generated by the policy based on the current state . The process continues until an end-of-sequence token is generated, completing the translation. The reference tokens in the target language are denoted by , where . The generated tokens are evaluated against to measure the quality of the translation. For , the state transition function defines the probability of transitioning from one state to another by appending a chosen token to the current translation, and the reward function assigns a real-valued reward to each transition , where and , based on the quality of the generated translation sequence. Explicitly, the reward function can be written as:
Sequence-level rewards are provided only once per translation and evaluate the entire output at once. Token-level rewards, on the other hand, give feedback for each generated token. The discount factor is used to weigh future rewards, with typically chosen in MT to ensure that all rewards are valued equally, allowing the optimization of the entire sequence of tokens in the translation rather than focusing on just the initial tokens. Finally, the goal is to maximize the expected cumulative reward over trajectories sampled from the . The objective function can be written as:
@@
To optimize the above objective, we can use policy gradient methods: REINFORCE Williams (1992); Ranzato et al. (2015), a vanilla policy gradient method, optimizes translation by sampling hypotheses , scoring them with a reward obtained from an MT metric , and updating the model to maximize expected rewards, as shown in Equation 2. Despite its simplicity, it often struggles with high variance and instability. Proximal Policy Optimization (PPO) Schulman et al. (2017) mitigates this by using a clipped surrogate objective (Equation 3) to keep policy updates stable within a margin , efficient and employs Generalized Advantage Estimation (GAE) Schulman et al. (2015) to compute advantages using rewards and value function . While PPO performs well across various tasks, simpler methods like REINFORCE can sometimes rival or surpass it Ahmadian et al. (2024). Both are evaluated in our experiments (§5.2).
@@ 3 Related Work
Incorporating human feedback as rewards and optimizing language models with RL methods effectively aligns them with human preferences Ouyang et al. (2022), often surpassing MLE. A notable example in translation tasks is Minimum Risk Training (MRT) Shen et al. (2016), which minimizes expected risk based on evaluation metrics to directly improve translation quality. Recent advances in NMT build on this idea by refining training with feedback from metrics or human evaluations, incorporating alignment techniques and RL methods Nguyen et al. (2017); Kreutzer et al. (2018b); Wu et al. (2018); Kiegeland and Kreutzer (2021); Xu et al. (2023); Agrawal et al. (2024); Zhu et al. (2024); He et al. (2024); Ramos et al. (2024). Despite these advancements, sentence-level feedback methods face persistent challenges such as sparse rewards, instability, and difficulty handling long sequences Wu et al. (2018). These issues hinder performance, generalization, and robust learning, even with multi-objective optimizations Wu et al. (2023); Jang et al. (2023). To address the limitations of sentence-level feedback, recent research has explored finer-grained rewards at the token level for tasks such as language model alignment Xia et al. (2024); Yoon et al. (2024); Guo et al. (2023); Cao et al. (2024), controllable text generation Li et al. (2024), query generation Ouyang et al. (2024), among others, but remain relatively underexplored in MT.
@@
Previous approaches to token-level reward modeling often relied on binary error markings generated by humans Kreutzer et al. (2020); Domingo et al. (2017) or simulated it by comparing model predictions with reference translations based on heuristic methods Petrushkov et al. (2018). While effective, these methods provide limited feedback due to their binary nature and require costly human annotation, making them less practical for scalable solutions. Other approaches have employed reward shaping techniques Ng et al. (1999); Wu et al. (2018); Goyal et al. (2019); Devidze et al. (2022), incorporating intermediate rewards along with BLEU Papineni et al. (2002) as the reward function. However, partial BLEU or token-level BLEU are less effective for fine-grained reward modeling, as they depend on exact -gram matching and fail to capture meaningful semantic differences and context. Consequently, these methods, while valuable, are limited in their granularity and fail to address the severity of errors introduced at the token level.
To address these limitations, in this work, we instead use token-level rewards obtained from state-of-the-art evaluation metrics that predict error spans and error severity levels to improve the learning and translation quality of MT systems.
@@ 4 Approach
In this section, we present our method for incorporating token-level rewards into RL training for machine translation (MT). We first detail how error assessments are mapped to token-level reward scores and then describe our adaptation of the REINFORCE and PPO learning objectives at the token level to enable learning from this fine-grained feedback.
@@
We use the evaluation metric, xCOMET as as it was shown to achieve the best correlation with human judgments and was the winning submission for the WMT23 Metrics Shared Task Freitag et al. (2023). The severity weights from xCOMET adhere to the MQM framework Lommel et al. (2014), which classifies translation issues into categories such as fluency, adequacy, grammar, and style. Each token within an error span is assigned the same severity weight (see Figure 1). We note that although the weights follow the MQM guidelines, they need to be further adjusted depending on the tasks to optimize the performance of token-level RL.
@@
MT as well as evaluation models often rely on subword-level tokenization methods, such as Byte-Pair Encoding (Gage, 1994, BPE) or SentencePiece Kudo and Richardson (2018), where a single word can be split into multiple tokens. For a detokenized hypothesis generated by the MT system, identifies detokenized error spans corresponding to the errors. To generate token-level reward assignments for the tokenized hypothesis, , we assign the same reward to all subword tokens that belong to the same parent word. This is done by identifying the boundaries of the original word and ensuring that all tokens within those boundaries share the same reward. In this way, token-level rewards remain robust and independent of the tokenization differences between the generation and the evaluation models.
@@
The policy ratio captures the change in policy for each token relative to the previous policy. The token-level advantage measures the improvement in performance over a baseline for each token and is used to scale updates based on expected reward improvements, as described in Schulman et al. (2015, 2017). To ensure stable training, we apply clipping to limit the extent of policy updates, preventing large, unstable shifts. This approach allows for more granular control over the model’s learning, ensuring that each token is generated in a way that maximizes task-specific objectives while maintaining stability in the policy updates. Clipping also helps mitigate length bias by preventing longer sequences from accumulating disproportionately high rewards.
@@ 5.1 Experimental Setup
We use three state-of-the-art models: a standard encoder-decoder MT model, NLLB111facebook/nllb-200-1.3B NLLB Team et al. (2022), and two LLM-based MT systems, Tower222Unbabel/TowerInstruct-Mistral-7B Alves et al. (2024) and Gemma333https://huggingface.co/google/gemma-2-9b-it Gemma Team et al. (2024). While NLLB and Tower are dedicated MT models optimized for translation tasks, Gemma is an LLM that exhibits strong multilingual capabilities. These models differ in both their architectures and pre-training methodologies. Each is pre-trained on diverse multilingual datasets, establishing them as robust baselines for investigating the effects of SFT and RL techniques.
@@
We use the following training datasets in our experiments: (1) The IWSLT2017 dataset Cettolo et al. (2017), with k examples for English-French (ENFR), supports rapid experimentation and frequent training iterations. (2) The WMT18 dataset Bojar et al. (2018) contains M examples for English-German (ENDE). We train NLLB with both datasets and the LLM-based models with (2). Training stops once rewards stabilize, so not all examples are used.
We evaluate NLLB models using the IWSLT17 ENFR and WMT18 ENDE test sets. To standardize comparison across MT systems (NLLB and LLM-based MT systems), we also evaluate all models on the WMT24 dataset Kocmi et al. (2024), addressing concerns about data contamination, as Tower training included the WMT18 test set.
@@
We assess translation quality using a comprehensive suite of well-established evaluation metrics. These include lexical reference-based metrics, such as BLEU Papineni et al. (2002) and ChrF Popović (2015); neural reference-based metrics, including COMET22 Rei et al. (2022), xCOMET Guerreiro et al. (2024), and BLEURT Sellam et al. (2020); and a neural reference-free metric, CometKiwi-23 Rei et al. (2023). Lexical metrics focus on word overlap and -gram matching, while neural metrics evaluate translations in terms of semantic coherence and contextual quality. Including a reference-free metric enables evaluation without reliance on predefined reference texts. This diverse set of metrics captures multiple dimensions of translation quality, including fluency, grammatical accuracy, semantic adequacy, and contextual relevance. By using various evaluation criteria, we reduce potential biases that may arise from aligning the reward model with a single evaluation metric, ensuring more robust and reliable conclusions about the impact of different approaches on translation quality.
We apply significance testing at a confidence threshold of . For segment-level metrics, such as COMET-22, we test at the segment level, but for corpus-level metrics, such as BLEU and ChrF, we apply bootstrapping with samples of size Koehn (2004). Performance clusters are formed based on statistically significant gaps, and final rankings are derived by averaging the cluster scores across all languages Colombo et al. (2022); Freitag et al. (2023).
In addition to automated metrics, we conduct human evaluations to gain deeper insights into translation performance. Furthermore, significance testing is performed to validate improvements observed across different models and approaches, ensuring the reliability of our findings.
@@
We utilize two reward models based on xCOMET444https://huggingface.co/Unbabel/XCOMET-XL. The MQM-derived reward signal, referred to as xCOMET-MQM, is generated from error span predictions that identify and classify translation errors by severity. Token-level reward signals are directly computed from these error spans, while sentence-level rewards are obtained as a weighted average of the token-level severity spans. We also use the standard sentence-level reward signal provided by xCOMET as a baseline for comparison.
@@
We finetune NLLB, Tower and Gemma models using MLE and RL methods detailed in §3 with the following configurations:
@@
CPO Xu et al. (2024): a state-of-the-art preference optimization learning method for MT, offering a more efficient variant of DPO Rafailov et al. (2023). We construct the preference dataset by generating multiple outputs from the MT model using the training datasets,555We generate 16 samples with the value of top_p top_k set to 0.9 and 50 respectively. and then induce preferences using the xCOMET metric, comparing these outputs to human-written references.
@@
We use the same hyperparameter settings for NLLB, Tower, and Gemma. We use HuggingFace’s Transformers library Wolf et al. (2020) and the Transformers Reinforcement Learning (TRL) library 666https://github.com/huggingface/trl to facilitate RL training. We perform MLE training with Adam Kingma and Ba (2014) as the optimization algorithm, learning rate decay starting from and early stopping. We use PPO with a learning rate of , set as , trajectory limit set as . Mini-batch updates are performed with a batch size of over PPO epochs. The translation prompt for LLM-based MT systems is shown below:

### v3

@@ 1 Introduction
Neural machine translation (NMT) Kalchbrenner2013RecurrentCT; sutskever2014sequence; cho2014properties, a leading approach within MT, leverages neural networks to automate language translation and has driven significant improvements in translation quality. However, most NMT systems are predominantly trained using maximum likelihood estimation (MLE). MLE-based training focuses on maximizing the probability of next-word predictions given a partial reference. This often leads to a critical problem known as exposure bias – the model uses ground-truth prefix tokens during training, but during inference it relies on its previous predictions BENGIO2015; RANZATO2016; WISEMAN2016. This can cause errors to propagate through the generated sequence, severely degrading the translation quality. Furthermore, it tends to produce translations that lack global coherence and adequacy as the model does not sufficiently consider the context of entire sentences or the overarching meaning. This has spurred interest in using alternative approaches that leverage RL methods for training NMT systems.
@@
RL-based approaches use explicit reward models to evaluate the outputs generated by the NMT system, assigning scores to generated hypotheses to guide the learning process. However, most prior research RANZATO2016; wu2016google; Bahdanau2016; nguyen-etal-2017-reinforcement; wuseq2017; kreutzer-etal-2018-neural; kreutzer-etal-2018-reliability; kiegeland-kreutzer-2021-revisiting predominantly relies on sentence-level feedback and often struggles with reward sparsity, particularly for long-form text generation: sentence-level rewards fail to capture specific issues within a translation, making it difficult for the model to learn from negative reward signals. As shown in Figure 1, two translations corresponding to different source texts of varying length receive the same sentence-level quality score of 70, yet differ significantly in the nature and impact of the errors: The first translation has several minor errors scattered throughout the text, while the latter has major errors that could potentially hinder the understanding of the original content. This suggests that learning can be more effective if feedback is provided at a fine-grained level, including precise identification of the nature of errors.
@@
Recent advancements in automated MT evaluation metrics that generate fine-grained error span predictions, such as xCOMET xcomet, MetricX juraska-etal-2023-metricx, AutoMQM fernandes-etal-2023-devil, EAPrompt (lu-etal-2024-error), MaTESE perrella-etal-2022-matese, and BARTScore++ (lu-etal-2023-toward) have shown promise in improving alignment with human translation quality judgments. These metrics directly predict token-level error severity (no error, minor, major, or critical) and optionally provide sentence-level quality assessments or prompt large language models to identify error types (e.g., mistranslation, omission) and severities based on the Multidimensional Quality Metrics (MQM) framework (lommel-mqm).
@@
Despite the potential of severity-based metrics to improve translation quality, their application in MT training via RL methods remains relatively underexplored, since it presents several challenges: (i) the feedback, albeit informative and frequent, can be noisy, and (ii) determining the appropriate reward assignments for different severity levels to ensure effective and stable learning is not straightforward. In this regard, our research aims to answer the following questions:
@@
Can fine-grained MT metrics be effectively used to provide accurate, detailed, human-aligned feedback to reduce reward sparsity?
@@
By integrating fine-grained reward signals into NMT training, we demonstrate significant improvements in translation quality and overcome the challenges of exposure bias, reward sparsity, and instability of RL training, paving the way for more reliable and accurate MT systems.
@@ 2 Background
In the context of MT, we can model the translation process as a Markov Decision Process (MDP) PUTERMAN1990, defined by the tuple with a finite vocabulary . The state space consists of all possible sequences of tokens up to the current time step, which includes the input sequence in the source language, as well as the target language tokens generated so far. Initially, the state corresponds to the input sentence in the source language, , where each token . At each time step , the state represents the sequence of tokens generated up to that point, which can be expressed as:
The agent selects an action , which is a token generated by the policy based on the current state . The process continues until an end-of-sequence token is generated, completing the translation. The reference tokens in the target language are denoted by , where . The generated tokens are evaluated against to measure the quality of the translation. For , the state transition function defines the probability of transitioning from one state to another by appending a chosen token to the current translation, and the reward function assigns a real-valued reward to each transition , where and , based on the quality of the generated translation sequence. Conceptually, the reward function is defined as a mapping from a hypothesis to a score, i.e., . In practice, many MT metrics additionally condition on the source and/or the reference, which we make explicit as . Formally, the reward function can be written as:
Sentence-level rewards are provided only once per translation and evaluate the entire output at once. Token-level rewards, on the other hand, give feedback for each generated token.The discount factor is used to weigh future rewards, with typically chosen in MT to ensure that all rewards are valued equally, allowing the optimization of the entire sequence of tokens in the translation rather than focusing on just the initial tokens. Finally, the goal is to maximize the expected cumulative reward over trajectories sampled from the . The objective function can be written as:
@@
To optimize the above objective, we can use policy gradient methods: REINFORCE WILLIAMS1992; RANZATO2016, a vanilla policy gradient method, optimizes translation by sampling hypotheses , scoring them with a reward obtained from an MT metric , and updating the model to maximize expected rewards, as shown in Equation 2. Despite its simplicity, it often struggles with high variance and instability. Proximal Policy Optimization (PPO) PPO mitigates this by using a clipped surrogate objective (Equation 3) to keep policy updates stable within a margin , efficient and employs Generalized Advantage Estimation (GAE) schulman2018highdimensionalcontinuouscontrolusing to compute advantages using rewards and value function . While PPO performs well across various tasks, simpler methods like REINFORCE can sometimes rival or surpass it ahmadian2024basics. Both are evaluated in our experiments (§5.2).
@@ 3 Related Work
Incorporating human feedback as rewards and optimizing language models with RL methods effectively aligns them with human preferences ouyang2022training, often surpassing MLE. A notable example in translation tasks is Minimum Risk Training (MRT) shen2016minimumrisktrainingneural, which minimizes expected risk based on evaluation metrics to directly improve translation quality. Recent advances in NMT build on this idea by refining training with feedback from metrics or human evaluations, incorporating alignment techniques and RL methods nguyen-etal-2017-reinforcement; kreutzer-etal-2018-reliability; wu-etal-2018-study; kiegeland-kreutzer-2021-revisiting; almapref; agrawal-etal-2024-modeling; Zhu2024; he-etal-2024-improving; ramos-aligning. Despite these advancements, sentence-level feedback methods face persistent challenges such as sparse rewards, instability, and difficulty handling long sequences wu-etal-2018-study. These issues hinder performance, generalization, and robust learning, even with multi-objective optimizations wu2023finegrained; jang2023personalizedsoupspersonalizedlarge. To address the limitations of sentence-level feedback, recent research has explored finer-grained rewards at the token level for tasks such as language model alignment xia-etal-2024-inverse; yoon-etal-2024-tlcr; 2023arXiv231104072G; cao-etal-2024-enhancing, controllable text generation li-etal-2024-reinforcement, query generation 2024arXiv241100722O, among others, but remain relatively underexplored in MT.
@@
Previous approaches to token-level reward modeling often relied on binary error markings generated by humans kreutzer-etal-2020-correct; domingo2017segment or simulated it by comparing model predictions with reference translations based on heuristic methods petrushkov-etal-2018-learning. While effective, these methods provide limited feedback due to their binary nature and require costly human annotation, making them less practical for scalable solutions. Other approaches have employed reward shaping techniques ng1999policy; wu-etal-2018-study; RS_GOYAL; RS_RATI, incorporating intermediate rewards along with BLEU bleu as the reward function. However, partial BLEU or token-level BLEU are less effective for fine-grained reward modeling, as they depend on exact -gram matching and fail to capture meaningful semantic differences and context. Consequently, these methods, while valuable, are limited in their granularity and fail to address the severity of errors introduced at the token level.
@@ 4 Approach
In this section, we present our method for incorporating token-level rewards into RL training for machine translation (MT). To address the limitations of prior approaches, such as binary feedback or coarse sentence-level scores, we use token-level rewards derived from state-of-the-art evaluation metrics that predict error spans and severity levels. These fine-grained signals are then used to guide learning through adaptations of REINFORCE and PPO objectives at the token level, enabling more effective and stable training of MT systems.
@@
We use the evaluation metric, xCOMET as as it was shown to achieve the best correlation with human judgments and was the winning submission for the WMT23 Metrics Shared Task freitag-etal-2023-results. The severity weights from xCOMET adhere to the MQM framework lommel-mqm, which classifies translation issues into categories such as fluency, adequacy, grammar, and style. Each token within an error span is assigned the same severity weight (see Figure 1). We note that although the weights follow the MQM guidelines, they need to be further adjusted depending on the tasks to optimize the performance of token-level RL.
@@
MT systems and evaluation models typically use subword-level tokenization methods such as Byte-Pair Encoding (BPE, BPE) or SentencePiece kudo-richardson-2018-sentencepiece, where words can be split into multiple subword tokens and token boundaries may not align with natural word boundaries. Given a detokenized hypothesis from the MT system, our evaluation model produces error spans defined at the character level. To assign rewards at the token level for the tokenized hypothesis , we first re-tokenize the detokenized hypothesis using the same tokenizer applied during model training. This allows us to obtain precise character offsets for each subword token. We then align tokens to the character-level error spans by checking for overlap: any token whose character span overlaps with an error span inherits the corresponding error severity. This alignment avoids relying on explicit word boundaries or whitespace segmentation, making the reward assignment robust to different tokenization schemes – including those that generate cross-word subword units – and applicable across languages with or without explicit word boundaries. By grounding token-level rewards in character-level overlap rather than word-based grouping, our method ensures consistency and generalizability across tokenization models and languages. Finally, if a token overlaps multiple spans, it is assigned the worst severity – critical major minor correct – avoiding averaging or length-weighting to remain tokenizer-agnostic. Formally, for a token with overlaps , if is empty, . This severity is then mapped to its numeric reward (Table 4). The full algorithm is provided in Appendix A.
@@
The policy ratio captures the change in policy for each token relative to the previous policy. The token-level advantage is estimated using Generalized Advantage Estimation (GAE) (schulman2018highdimensionalcontinuouscontrolusing), which balances bias and variance by mixing temporal-difference (TD) errors across multiple steps. The advantage at time step is computed as:
where is the temporal-difference (TD) error, is the GAE parameter, is the reward at time step , and is a learned value function that estimates the expected return from the state defined by the input and the generated prefix . As explained earlier, we consider for our use case, and therefore omit the discount factor for notational simplicity. To ensure stable training, we apply clipping to limit the extent of policy updates, preventing large, unstable shifts. This approach allows for more granular control over the model’s learning, ensuring that each token is generated in a way that maximizes task-specific objectives while maintaining stability in the policy updates. Clipping also helps mitigate length bias by preventing longer sequences from accumulating disproportionately high rewards.
@@ 5.1 Experimental Setup
We use three state-of-the-art models: a standard encoder-decoder MT model, NLLB flores, and two LLM-based MT systems, Tower tower and Gemma gemmateam2024gemma2improvingopen. While NLLB and Tower are dedicated MT models optimized for translation tasks, Gemma is an LLM that exhibits strong multilingual capabilities. These models differ in both their architectures and pre-training methodologies. Each is pre-trained on diverse multilingual datasets, establishing them as robust baselines for investigating the effects of SFT and RL techniques.
@@
We use the following training datasets in our experiments: (1) The IWSLT2017 dataset iwslt2017, with k examples for English-French (ENFR), supports rapid experimentation and frequent training iterations. (2) The WMT18 dataset WMT18 contains M examples for English-German (ENDE). We train NLLB with both datasets and the LLM-based models with (2). Training stops once rewards stabilize, so not all examples are used. We evaluate NLLB models using their respective test splits: IWSLT17 (ENFR) and WMT18 (ENDE). To standardize comparison across MT systems (NLLB and LLM-based MT systems), we also evaluate all models on the WMT24 dataset WMT24, addressing concerns about data contamination, as Tower training included the WMT18 test set.
@@
We assess translation quality using a comprehensive suite of well-established evaluation metrics. These include lexical reference-based metrics, such as BLEU bleu and ChrF chrf; neural reference-based metrics, including COMET22 comet22, xCOMET xcomet, and BLEURT bleurt; and a neural reference-free metric, CometKiwi-23 cometkiwi. Lexical metrics focus on word overlap and -gram matching, while neural metrics evaluate translations in terms of semantic coherence and contextual quality. Including a reference-free metric enables evaluation without reliance on predefined reference texts. This diverse set of metrics captures multiple dimensions of translation quality, including fluency, grammatical accuracy, semantic adequacy, and contextual relevance. By using various evaluation criteria, we reduce potential biases that may arise from aligning the reward model with a single evaluation metric, ensuring more robust and reliable conclusions about the impact of different approaches on translation quality.
We apply significance testing at a confidence threshold of . For segment-level metrics, such as COMET-22, we test at the segment level, but for corpus-level metrics, such as BLEU and ChrF, we apply bootstrapping with samples of size koehn-2004-statistical. Performance clusters are formed based on statistically significant gaps, and final rankings are derived by averaging the cluster scores across all languages colombo2022bestsystemsnewperspectives; freitag-etal-2023-results. In addition to automated metrics, we conduct human evaluations with two professional annotators, reporting inter-annotator agreement (Pearson’s and Spearman’s ) and confidence intervals across length buckets to assess the reliability of model differences.
@@
We utilize two reward models based on xCOMET. The MQM-derived reward signal, referred to as xCOMET-MQM, is generated from error span predictions that identify and classify translation errors by severity. Token-level reward signals are directly computed from these error spans, while sentence-level rewards are obtained as a weighted average of the token-level severity spans. We also use the standard sentence-level reward signal provided by xCOMET as a baseline for comparison.
@@
We finetune NLLB, Tower and Gemma models using MLE and RL methods detailed in Section 3 with the following configurations:
@@
CPO CPO: a state-of-the-art preference optimization learning method for MT, offering a more efficient variant of DPO DPO. We construct the preference dataset by generating multiple outputs from the MT model using the training datasets,111We generate 16 samples with the value of top_p top_k set to 0.9 and 50 respectively. and then induce preferences using the xCOMET metric, comparing these outputs to human-written references.
@@
We use the same hyperparameter settings for NLLB, Tower, and Gemma. We use HuggingFace’s Transformers library wolf-etal-2020-transformers and the Transformers Reinforcement Learning (TRL) library to facilitate RL training. We perform MLE training with Adam kingma2017adam as the optimization algorithm, learning rate decay starting from and early stopping. We use PPO with a learning rate of , set as , trajectory limit set as . Mini-batch updates are performed with a batch size of over PPO epochs. The translation prompt for LLM-based MT systems is shown in Table 1.